the sum of 4 numbersis 726. the average of the first three numbers is 149. what is the forth number ?
Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52863)

.
the sum of 4 numbersis 726. the average of the first three numbers is 149. what is the forth number ?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

It is 726 - 3*149.


Answer by stanbon(75887)

the sum of 4 numbersis 726. the average of the first three numbers is 149. what is the forth number ?
------
Sum of the 1st 3 numbers is 3*149 = 437
-----
4th number = sum of 4 numbers - sum of 3 numbers = 726-437 = 289
-----------
Cheers,
Stan H.
-------------